# Emergence of the Primacy Effect in Structured State-Space Models

### v1

@@
Emergence of the Primacy Effect in Structured State-Space Models
Takashi Morita Academy of Emerging Sciences, Chubu University
1 Abstract
Human and animal memory for sequentially presented items is well-documented to be more accurate for those at the beginning and end of a sequence, phenomena known as the primacy and recency effects, respectively. By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time. Accordingly, ANNs are expected to show the recency effect but not the primacy effect. Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture, called structured state-space models, exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
@@ 2 Introduction
Human and animal memory for sequentially presented items is well-documented to be more accurate for those appearing at the beginning and end of a sequence—phenomena known as the primacy and recency effects, respectively (Ebbinghaus,, 1913; Murdock,, 1962; Glanzer and Cunitz,, 1966). For example, when a sequence of random integers such as $49,75,\dots,5,38$ is presented in that order, the initial ($49,75$) and final ($5,38$) items are more likely to be recalled accurately at the end of the presentation.
@@ 3.1 Recurrent Neural Networks
The traditional ANN-based approach to time-series processing relies on recurrent neural networks (RNNs; Elman,, 1990). RNNs sequentially process inputs, updating their internal state (represented by a real-valued vector) at each time step based on both the current input and the previous state. The output is then generated through a feedforward transformation of this internal state at each step. In other words, RNNs map input time series to target output signals through the latent dynamics of their internal state.
Theoretically, RNNs possess Turing-complete computational power, meaning they can simulate an arbitrary computational model in an “idealized” setting (Siegelmann and Sontag,, 1992, 1994, 1995; Siegelmann,, 1996, 1999). This theoretical capability has driven widespread applications of RNNs in domains such as language modeling (Sundermeyer et al.,, 2012; Graves,, 2013), speech synthesis (Kalchbrenner et al.,, 2018), and weather nowcasting (Shi et al.,, 2015).
@@ 3.2 Structured State-Space Models
where $P_{n}$ denotes the basis polynomial of degree $n$ and $h_{n}(t)$ are the optimal coefficients for the approximation at time $t$111 The optimal coefficients can be determined as $h_{n}(t):=\langle x|_{\leq t},P_{n}\rangle=\int x|_{\leq t}(s)P_{n}(s)d\mu^{(t)}(s)$ when $\{P_{n}\}_{n=0}^{N-1}$ form an orthogonal basis with respect to the time-dependent measure $d\mu^{(t)}(\cdot)$. Then, these coefficients offer a finite- and constant-dimensional representation of the input signal up to time $t$. The framework can be naturally extended to multi-channel signals by performing channel-wise approximations in parallel, yielding $h_{n}^{(m)}(t)$ for each channel $m$.
@@
where $\mathbf{h}(t):=(h_{1}(t),\dots,h_{N}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t):=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
@@
where $\Delta t:=t_{j+1}-t_{j}$ ($\forall j=1,\dots,L-1$) defines the time-step size. This time-step parameter is treated as learnable (for each input channel), allowing the model to automatically adjust the time scale of its state-space dynamics to align with that of the input signal.
@@
It is also important to remember that discretization methods are generally designed under the assumption of sufficiently small $\Delta t$, for effectively approximating the limit $\Delta t\to 0$. Consequently, setting $\Delta t$ too large introduces discrepancies between the continuous and discretized dynamics (as illustrated by the contrast between the left vs. right panels in Fig. 1).
@@ 3.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at the current time $t$ while maintaining full access to the most recent $M$ observations, $x_{t-M-1},\dots,x_{t}$. At each time step, they “forget” or discard the oldest observation ($x_{t-M}$) as new data enter the storage buffer.
By definition, finite-queue/Markov models do not exhibit the primacy effect, as the earliest observations are inevitably lost once they exceed the finite storage window. Instead, these models align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). Given this fundamental limitation, finite-queue/Markov models fall outside the scope of the present study.
@@
4 Methods
4.1 Task
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).444 The most well-adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in a theoretically more demanding task requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
@@ 4.1 Task
To build test sequences, the held-out study items were randomly ordered, and queries were first generated by first shuffling and then cyclically shifting them (e.g., $(2,8,11,29)\mapsto\{(2,8,11,29),(8,11,29,2),(11,29,2,8),(29,2,8,11)\}$). This design ensured that each study item was queried in all $L$ possible positions. Finally, either the even- or odd-indexed query positions were replaced with random distractors, resulting in a total of $1024\times L\times 2$ test sequences per trial.
@@ 4.2 Models
The models used for the binary memory verification task comprised three layers, as illustrated in Fig. 2. First, the input integers were embedded into 256-dimensional real-valued vectors. These embeddings were shared between study items and query tokens. The resulting sequence of vectors was then processed by the SSM/RNN, whose outputs were linearly projected onto binary logits to determine whether each query token was present in the study items.
@@
For comparison, a single-layer long short-term memory (LSTM) network was also evaluated (Hochreiter and Schmidhuber,, 1997). LSTM has been the goldstandard RNN architecture for various time-series processing tasks, including language modeling (Sundermeyer et al.,, 2012; Graves,, 2013). The dimensionality of both hidden and cell states of was set to 256.
@@ 5.1 Emergence of the Primacy Effect
Figure 3: Accuracy of the binary memory verification task. Each cell in the heatmap represents the accuracy of study items that were first observed at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the S4 model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ ($0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
@@ 5.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ is a key parameter governing the memory capacity of the SSM, although the primacy effect emerged across different initial values (Fig. 3). To further investigate its role, the learning trajectories of $\Delta t$ over training iterations were analyzed. The results indicated that as training progressed, a specific range of step sizes became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig 4).
@@ 6.1 Summary of the Findings
The present study demonstrated that the SSM exhibits the primacy effect in memorization. When performing the binary memory verification task, which parallels paradigms used to investigate memory capacities in humans and animals (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985), the model showed the highest accuracy for study items presented at the beginning of the sequence. Moreover, memorized information was not retrievable immediately after the presentation of the study items. These findings are novel and counterintuitive, as they challenge the theoretical formulation of the SSM, which assumes an exponentially decaying measure (for Legendre and Laguerre bases; Fig 1; Gu et al.,, 2023).
@@
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored.666 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
@@ 6.2 Comparison with Data/Task-Oriented Primacy Effects
Several prior studies have reported a primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) examined positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.777 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: ”) consistently corresponded to their order of presentation. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels.
@@
Since the Transformer architecture is inherently position-agnostic—lacking an intrinsic ordering mechanism apart from external positional encodings (Vaswani et al.,, 2017)—the primacy effects observed in these studies must stem from the structure of the training data or task design. Wang et al., argued that LLMs inherit cognitive biases from human-generated linguistic data. Xiao et al., suggested that the nature of the language modeling task itself encourages prioritization of initial tokens, as they are repeatedly used as inputs for autoregressive predictions, reinforcing attention allocation to them.

### v2

@@
Abstract
Human and animal memory for sequentially presented items is well-documented to be more accurate for those at the beginning and end of the sequence, phenomena known as the primacy and recency effects, respectively. By contrast, artificial neural network (ANN) models are typically designed with a memory that decays monotonically over time. Accordingly, ANNs are expected to show the recency effect but not the primacy effect. Contrary to this theoretical expectation, however, the present study reveals a counterintuitive finding: a recently developed ANN architecture, called structured state-space models, exhibits the primacy effect when trained and evaluated on a synthetic task that mirrors psychological memory experiments. Given that this model was originally designed for recovering neuronal activity patterns observed in biological brains, this result provides a novel perspective on the psychological primacy effect while also posing a non-trivial puzzle for the current theories in machine learning.
@@ 1 Introduction
Human and animal memory for sequentially presented items is well-documented to be more accurate for those appearing at the beginning and end of the sequence—phenomena known as the primacy and recency effects, respectively (Ebbinghaus,, 1913; Murdock,, 1962; Glanzer and Cunitz,, 1966). For example, when a sequence of random integers such as $49,75,\dots,5,38$ is presented in that order, the initial ($49,75$) and final ($5,38$) items are more likely to be recalled accurately at the end of the presentation.
@@ 2.1 Recurrent Neural Networks
The traditional ANN-based approach to time-series processing relies on recurrent neural networks (RNNs; Elman,, 1990). RNNs sequentially process inputs, updating their internal state (represented by a real-valued vector) at each time step based on both the current input and the previous state. The output is then generated through a feedforward transformation of this internal state at each step. In other words, RNNs map input time series to target output sequences through the latent dynamics of their internal state.
Theoretically, RNNs possess Turing-complete computational power, meaning they can simulate an arbitrary computational model in an “idealized” setting (Siegelmann and Sontag,, 1992, 1995; Siegelmann,, 1999). This theoretical capability has driven widespread applications of RNNs in domains such as language modeling (Sundermeyer et al.,, 2012; Graves,, 2013), speech synthesis (Kalchbrenner et al.,, 2018), and weather nowcasting (Shi et al.,, 2015).
@@ 2.2 Structured State-Space Models
where $P_{n}$ denotes the basis polynomial of degree $n$ and $h_{n}(t)$ are the optimal coefficients for the approximation at time $t$111 The optimal coefficients can be determined as $h_{n}(t)=\langle x|_{\leq t},P_{n}\rangle=\int x|_{\leq t}(s)P_{n}(s)d\mu^{(t)}(s)$ when $\{P_{n}\}_{n=0}^{N-1}$ form an orthogonal basis with respect to the time-dependent measure $d\mu^{(t)}(\cdot)$. Then, these coefficients offer a finite- and constant-dimensional representation of the input signal up to time $t$. The framework can be naturally extended to multi-channel signals by performing channel-wise approximations in parallel, yielding $h_{n}^{(m)}(t)$ for each channel $m$.
@@
where $\mathbf{h}(t):=(h_{1}(t),\dots,h_{N}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)=C\mathbf{h}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t)=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
@@
where $\Delta t:=t_{j+1}-t_{j}$ ($\forall j=1,\dots,L-1$) defines the time-step size. This time-step parameter is treated as learnable, allowing the model to automatically adjust the time scale of its state-space dynamics to align with that of the input signal. Moreover, in multi-channel settings, the model can represent multi-scale dynamics by assigning distinct $\Delta t^{(m)}$ values to each channel $m$.
@@
It is also important to note that discretization methods are generally designed under the assumption of sufficiently small $\Delta t$, for effectively approximating the limit $\Delta t\to 0$. Consequently, setting $\Delta t$ too large introduces discrepancies between the continuous and discretized dynamics (as illustrated by the contrast between the left vs. right panels in Fig. 1).
@@ 2.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at the current discrete time $j$ while maintaining full access to the most recent $M$ observations, $x_{j-M+1},\dots,x_{j}$. At each time step, they “forget” or discard the oldest observation ($x_{j-M}$) as new data enter the storage buffer.
By definition, finite-queue/Markov models do not exhibit the primacy effect, in a sense that the earliest observations are inevitably lost once they exceed the finite storage window. Instead, these models align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). Given this fundamental limitation, finite-queue/Markov models fall outside the scope of the present study.
3 Methods
3.1 Task
@@
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).444 The most widely adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be technically formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in theoretically more demanding tasks requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
@@
To build test sequences, the held-out study items were randomly ordered, and queries were generated by first shuffling and then cyclically shifting them (e.g., $(2,8,11,29)\mapsto\{(2,8,11,29),(8,11,29,2),(11,29,2,8),(29,2,8,11)\}$). This design ensured that each study item was queried in all $L$ possible positions. Finally, either the even- or odd-indexed query positions were replaced with random distractors, resulting in a total of $1024\times L\times 2$ test sequences per trial.
@@ 3.2 Models
The models used for the binary memory verification task comprised three layers, as illustrated in Fig. 2. In the first layer, the input integers were embedded into 256-dimensional real-valued vectors. These embeddings were shared between study items and query tokens. The resulting sequence of vectors was then processed by the SSM/RNN, whose outputs were linearly projected onto binary logits to determine whether each query token was present in the study items.
@@
For comparison, a single-layer long short-term memory (LSTM) network was also evaluated (Hochreiter and Schmidhuber,, 1997). LSTM has been the goldstandard RNN architecture for various time-series processing tasks, including language modeling (Sundermeyer et al.,, 2012; Graves,, 2013). The dimensionality of both hidden and cell states was set to 256.
@@
Figure 3: Accuracy of the binary memory verification task. Each cell in the heatmap represents the accuracy for study items that were first observed at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the S4 model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ (i.e., $0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
4 Results
4.1 Emergence of the Primacy Effect
@@ 4.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ is a key parameter governing the memory capacity of the SSM, although the primacy effect emerged across different initial values (Fig. 3). To further investigate its role, the learning trajectories of $\Delta t$ over training iterations were analyzed. The results indicated that as training progressed, a specific range of step sizes ($\Delta t\leq 0.03$) became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig 4).
@@ 5.1 Summary of the Findings
The present study demonstrated that the SSM exhibits the primacy effect in memorization. When performing the binary memory verification task, which parallels paradigms used to investigate memory capacity in humans and animals (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985), the model showed the highest accuracy for study items presented at the beginning of the sequence. Moreover, memorized information was not retrievable immediately after the presentation of the study items. These findings are novel and counterintuitive, as they challenge the theoretical formulation of the SSM, which assumes an exponentially decaying measure (for Legendre and Laguerre bases; Fig 1; Gu et al.,, 2023).
@@
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit.666 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\Delta t\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
@@ 5.2 Comparison with Data/Task-Oriented Primacy Effects
Several prior studies have reported a primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) examined positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.777 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: card_arrival”) consistently corresponded to their order of presentation; that is, “Label $j$: xxx” was never listed before “Label $j^{\prime}$: yyy” if $j^{\prime}<j$, even though the label contents (“xxx” and “yyy”) were disassociated from the label indices and ordered randomly. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels.
@@
Since the Transformer architecture is inherently position-agnostic—lacking an intrinsic ordering mechanism apart from external positional encodings (Vaswani et al.,, 2017)—the primacy effects observed in these studies must stem from the statistical properties of the training data or task design. Wang et al., argued that LLMs inherit cognitive biases from human-generated linguistic data. Xiao et al., suggested that the nature of the language modeling task itself encourages prioritization of initial tokens, as they are repeatedly used as inputs for autoregressive predictions, reinforcing attention allocation to them.